OpenStudy (anonymous):

Solve the following equations. 4x - 8 = 2(2x - 3) -6x + 4 = -8x - 12 anyone help me on this and explain it to me? -4x+4 - 2x = -2(3x -2)

OpenStudy (anonymous):

Distribute the 2 and -2 inside the parentheses first.

OpenStudy (anonymous):

ok i solved 2 of them. I just need help with -4x+4 -2x=-2(3x-2)

OpenStudy (anonymous):

Keep in mind that our end goal is to get x by itself on one side of the equation.\[-4x+4-2x=-2(3x-2)\]\[-6x+4=-6x+4\]I guess x = 0

OpenStudy (anonymous):

or 1 or anything

OpenStudy (anonymous):

x = all real numbers